# Input Reduction Enhanced LLM-based Program Repair

### v1

@@ 4.1. Models
Table 1 summarizes the 4 LLMs evaluated in this work. Our selection balances two practical factors: the ability to run an LLM locally and the availability of cloud endpoints with competitive token pricing. For the consumer‑GPU category, we select GLM‑4‑9B‑chat (GLM et al., 2024) and Qwen2.5‑Coder‑7B (Hui et al., 2024). Both are fully open-source, fit comfortably on a single 24 GB consumer GPU, and therefore incur no api costs. To represent low‑cost hosted offerings, we add Qwen2.5‑Plus (Hui et al., 2024) and DeepSeek‑V3 (Liu et al., 2024). Qwen2.5‑Plus is a closed‑weight variant of the Qwen2.5 family (the provider does not disclose an exact parameter count) and is priced at \$0.11 per million input tokens and \$0.27 per million output tokens. DeepSeek‑V3 is larger (670B parameters, 37B active during decoding) yet still affordable at \$0.27 per million input tokens and \$1.11 per million output tokens. All select LLMs are the standard chat versions rather than the more expensive thinking variants (such as DeepSeek-R1), ensuring a fair and comparable inference budget. This selection enables us to examine how input reduction behaves on both locally deployed LLMs and cost-efficient cloud LLMs.
Table 1. LLMs used in this study.
Model Params (B) Cutoff \$ Cost per 1M tokens
Input Output
GLM‑4‑9B‑chat 9 Oct. 2023 / /
Qwen2.5‑Coder‑7B 7.6 Mar. 2024 / /
Qwen2.5‑Plus N/A Mar. 2024 0.11 0.27
DeepSeek‑V3 671 Jun. 2024 0.27 1.11
@@ 4.2. Benchmark
A fair study of input reduction for automated program repair (APR) requires two features that existing datasets lack: (1) long failure‑inducing test inputs and (2) low risk of training leakage. Widely used suites such as Defects4J (Just et al., 2014), Human‑EvalFix (Muennighoff et al., [n. d.]), and TutorCode (Yang et al., 2024) include only tiny tests, usually a few hundred characters, so they do not reveal how well an APR pipeline copes when the failing input grows to tens of kilobytes. Meanwhile, most of those benchmarks were released years ago and are drawn from popular open‑source projects that large language models have almost certainly seen, which can overstate repair accuracy. To our knowledge, no existing benchmark simultaneously offers long failure-inducing tests and low leakage risk; therefore, we need to build a new benchmark to fill this gap.
To meet the two requirements, the data source must publish every test file, including the largest hidden cases, and it must appear after the training cut‑off dates of selected LLMs. Codeforces does not satisfy the first point because it shares only small samples. AtCoder, by contrast, released the full test archives for every Beginner Contest (ABC) up to ABC 377, providing us with large inputs along with an official oracle. Therefore, we select all the satisfied tasks from ABC contests numbered 361 to 377, a span entirely after the knowledge cut‑offs of the 4 LLMs we evaluate. From each contest, we retain tasks whose largest official test file is at least 4 KB and whose difficulty level is between C and F, ensuring that the tasks remain solvable for LLMs. Table 2 lists the chosen 20 tasks.
Table 2. Composition of LFTBench.
Difficulty Task IDs # Tasks # Buggy Codes
C
361C, 366C, 368C, 375C,
376C, 377C
6 60
D
362D, 364D, 365D, 367D,
369D, 370D, 371D, 376D
8 80
E&F
363E, 372E, 373E, 374E,
376E, 377F
6 60
Total 20 200
Therefore, the benchmark contains 20 tasks: 6 C, 8 D, and 6 E/F. For each task, we collect the ten most recent C++ submissions that failed on a large test before July 1, 2025, resulting in a total of 200 bugs. The median failing‑input size is over 1 MB, and the largest single file exceeds 8 MB, which is large enough to trigger the “lost‑in‑the‑middle” effect.
@@ 4.3. Metrics
A repair attempt succeeds when at least one candidate patch passes the complete test suite (Eq. 3). Let $C_{b}(k)$ denote the event that bug $b$ is fixed by any of the first $k$ independent samples. Following common practice, we measure
[…]
and report pass@1, pass@5, and pass@10. These three cut‑offs align with how developers inspect automated suggestions in practice. Kochhar et al. (Kochhar et al., 2016) observe that most developers stop using a debugging tool if it does not help within the first five attempts, and Noller et al. (Noller et al., 2022) find that few users review more than ten ranked patches. The same thresholds are widely adopted in recent APR work (Huq et al., 2022; Fu et al., 2022; Luo et al., 2025; Yang et al., 2025c). Thus pass@1 reflects a one‑shot setting, while pass@5 and pass@10 model realistic batch sizes that can be screened offline without taxing developer patience.
@@ 4.4. Hyperparameters
Table 3. Key hyper‑parameters in ReduceFix.
@@
Reducer generation LLM backend Qwen‑Plus
@@
Wall‑clock limit 60s per reduction
Repair inference # Samples per bug (pass@k) $k=1,5,10$
@@ 5.1. Research Questions
  • RQ-1: How reliable are LLM-generated reducers at shrinking failure-inducing inputs? We assess the ReduceFix reduction phase through two metrics: the success rate and the compression ratio. Both metrics are evaluated against the unreduced original failure-inducing input and a purely LLM-based input reduction approach.
  • RQ-2: Does supplying the reduced counterexample improve LLM-based repair? Using 4 LLMs, we test 3 prompting conditions: Baseline (no failure-inducing input), Origin Test (the full failure-inducing input), and Reduced Test (the reduced input produced by ReduceFix). We report pass@k for $k\in\{1,5,10\}$ to quantify any gain in repair accuracy.
@@
  • RQ-4: Can reduced‑input prompting complement existing LLM-based APR pipelines? We integrate the ReduceFix into the ChatRepair (Xia and Zhang, 2024) without modifying its logic. Comparing the augmented version against the original one measures whether input reduction provides a drop-in boost for third-party APR systems.
@@ 5.2. RQ-1: Effectiveness of LLM-generated Reducer
[Experimental Design]: We assess the reduction performance on LFTBench, which contains 200 buggy C++ submissions drawn from 20 Atcoder Beginner Contest tasks. For each task, we prompt Qwen-Plus once, using a one-shot example (ABC330D) and the problem statement of the task to produce reducer.py (see Section 3.2). The script then runs a ddmin loop on the full failure-inducing test input to find a smaller input that still causes the output of the buggy and reference programs to differ. To test whether ddmin is necessary, we add a pure-LLM baseline, where the same LLM tries to generate a shorter test input directly. A reduction is counted as “Success” when it finishes on time without errors, and the failure is preserved. We report the success rate and compression ratio by difficulty and visualize the distribution of compression ratios using a violin plot.
Table 4. Reducer success rate: pure LLM one‑shot vs. ReduceFix (LLM + ddmin).
Difficulty Samples Success Rate
Pure LLM ReduceFix
C 60 0.63 1.00
D 80 0.16 0.96
E&F 60 0.48 0.88
Overall 200 0.40 0.95
[Experimental Results]: Using the prompt template in Listing 2, Qwen-Plus produced a syntactically valid reducer.py for all 200 buggy codes. Hence, all subsequent failures are due to the search phase rather than code-generation errors. Table 4 shows that the LLM-generated reducers succeed on 95% of the 200 buggy codes. All submissions of C-difficulty tasks are successfully reduced, while D-difficulty tasks and E & F-difficulty tasks reach 96% and 86% success, respectively. In every failure, the ddmin loop stopped without finding a smaller input, so the output equaled the original input. All the runs met the 60-second limitation.
Table 5. Compression-rate statistics of ReduceFix.
Difficulty Samples Mean (%) Median (%)
C 60 84.5 100.0
D 77 97.0 100.0
E&F 53 83.0 99.9
Overall 190 89.1 100.0
Statistics in Table 5 report the compression ratio, defined as Eq. 2, which means the percentage of bytes eliminated by the reducer. Under this definition, larger numbers indicate stronger compression: a value of 100% indicates that the input was reduced to almost nothing, while 0% means no reduction at all. The median removal rate is 100% across all difficulty levels, indicating that at least half of the test inputs that induce failures are nearly fully stripped yet still reproduce the bug. Averaged by difficulty, the reducer removes 83% of bytes on E & F-difficulty tasks, leaving 17% of the original data. Tasks in C-difficulty and D-difficulty group achieve even larger average removals of 84.5% and 97.0% respectively. The violin plot in Figure 3 confirms this pattern: most points cluster near the median (100%), while a long lower tail for hard tasks (i.e., difficulty group E and F) highlights a few cases with modest shrinkage that lower the mean.
Figure 3. Statistics of Compression Rate.
Table 4 shows what happens when the LLM is leveraged to generate a shorter test input in a one-shot setting, without the pipeline of ReduceFix. This pure LLM baseline succeeds on only 40% of the bugs overall and just 16% of the D-difficulty tasks, confirming that simple LLM-based one-shot test input reduction is ineffective and unreliable.
@@
Method Pure LLM LLM Reducer Times
@@
Table 6 compares the input/output token usage of two reduction strategies on LFTBench. The pure LLM approach, where the LLM tries to generate a shorter input in one pass, consumes about 5.7 million input tokens in total. The ReduceFix requires only 0.94 million input tokens for the same dataset. At current API prices, this comparison translates to \$0.632 versus \$0.017, 98% saving. Two factors drive this saving. First, the reducer script is generated on a per-task basis and then reused for every buggy submission of this task. Second, prompts in ReduceFix do not include the original failure-inducing input, which is often very large in LFTBench; they contain only the buggy code and the compact candidate input produced by ddmin. The pure LLM baseline must embed the entire long test case in the prompt, which greatly inflates the token cost.
[Failed Case Study]: For submission 62869553 of task ABC372E, the student program maintains, for every disjoint‑set root, an array that stores the twenty largest vertex identifiers encountered so far. A Type 1 query merges two components by copying only the ten largest numbers from the other component, then sorts the twenty numbers in descending order. A Type 2 query asks for the $k$-th largest vertex inside the root that contains $v$. When a large component is merged into a smaller one, some very large identifiers disappear; later requests with $k\geq 9$ therefore return incorrect answers.
@@
We resolved the issue by adding one simple guard. After the reducer renumbers the remaining vertices, it immediately reruns the interestingness test. If the failure no longer reproduces, the script rolls back to the original identifiers for that iteration instead of accepting the renumbered version. This small check, implemented in a few lines of Python, prevents the defect from being masked and allows the reduction process to continue correctly without any other modifications.
[RQ‑1] [Findings]: (1) ReduceFix successfully reduced on 95% of the 200 bugs. Reductions delete 80%–97% of inputs on average, leaving a minimal yet still failing input. (2) Pure LLM reduction fixes only 40% of bugs and costs 5.7M input tokens, significantly lower than ReduceFix’s 95%. [Insights]: (1) These results validate the core design of ReduceFix: leveraging LLM to generate a reducer and then driving a systematic ddmin search is both essential and efficient. (2) The success of this “LLM + classic search” pattern suggests that similar hybrids could improve other code-analysis tasks.
@@ 5.3. RQ2: Effectiveness of ReduceFix
[Objective]: We determine whether steering the repair model with the counter‑example reduced by ReduceFix improves the accuracy of generating a correct patch compared with (i) no failure-inducing input and (ii) the full failure-inducing input.
[Experimental Design]: For every one of the 4 selected LLMs in Table 1 (Qwen2.5‑Coder‑7B‑instruct, GLM4‑9B‑chat, Qwen‑Plus, DeepSeek‑V3), we test three prompting modes: Baseline (no test), Origin Test(full failure-inducing test), and Reduced Test. All other hyperparameters are fixed (temperature 0.8, $k\in{1,5,10}$, 10 candidate patches per bug), as shown in Table 4.4. Results are grouped by task difficulty.
Table 7. Pass@K (%) across 4 LLMs (deltas vs. Baseline; green = gain, red = drop).
Difficulty Baseline (No Test) Origin Test Reduced Test (ReduceFix)
pass@1 pass@5 pass@10 pass@1 pass@5 pass@10 pass@1 pass@5 pass@10
@@
C 5.5 17.1 23.3 3.7 ($-1.8\%$) 13.2 ($-3.9\%$) 20.0 ($-3.3\%$) 5.5 (+0.0%) 16.7 ($-0.4\%$) 25.0 ($+1.7\%$)
D 6.4 17.4 25.0 6.1 ($-0.2\%$) 17.4 (-0.0%) 23.8 ($-1.2\%$) 9.9 ($+3.5\%$) 26.0 ($+8.5\%$) 36.2 ($+11.2\%$)
E&F 1.3 5.9 10.0 1.8 ($+0.5\%$) 7.1 ($+1.2\%$) 11.7 ($+1.7\%$) 2.3 ($+1.0\%$) 8.5 ($+2.6\%$) 11.7 ($+1.7\%$)
Overall 4.6 13.9 20.0 4.1 ($-0.5\%$) 13.1 ($-0.8\%$) 19.0 ($-1.0\%$) 6.3 ($+1.7\%$) 17.9 ($+4.1\%$) 25.5 ($+5.5\%$)
GLM4-9B-chat
C 2.8 5.8 8.3 1.3 ($-1.5\%$) 3.8 ($-2.0\%$) 5.0 ($-3.3\%$) 2.2 ($-0.7\%$) 4.1 ($-1.7\%$) 5.0 ($-3.3\%$)
D 3.5 10.6 13.8 2.0 ($-1.5\%$) 7.3 ($-3.3\%$) 11.2 ($-2.5\%$) 5.8 ($+2.3\%$) 14.5 ($+3.9\%$) 20.0 ($+6.2\%$)
E&F 0.7 1.6 1.7 0.5 ($-0.2\%$) 1.5 ($-0.1\%$) 1.7 (+0.0%) 1.2 ($+0.5\%$) 1.7 ($+0.1\%$) 1.7 (+0.0%)
Overall 2.5 6.5 8.5 1.3 ($-1.1\%$) 4.5 ($-2.0\%$) 6.5 ($-2.0\%$) 3.3 ($+0.9\%$) 7.5 ($+1.1\%$) 10.0 ($+1.5\%$)
Qwen-Plus
C 37.3 60.3 68.3 31.8 ($-5.5\%$) 53.5 ($-6.8\%$) 63.3 ($-5.0\%$) 33.7 ($-3.7\%$) 53.5 ($-6.8\%$) 65.0 ($-3.3\%$)
D 40.1 55.6 60.0 39.9 ($-0.2\%$) 58.7 ($+3.1\%$) 61.3 ($+1.3\%$) 43.6 ($+3.5\%$) 58.7 ($+3.1\%$) 62.5 ($+2.5\%$)
E&F 22.5 39.7 48.3 24.0 ($+1.5\%$) 45.7 ($+6.0\%$) 51.7 ($+3.3\%$) 25.0 ($+2.5\%$) 46.5 ($+6.8\%$) 55.0 ($+6.7\%$)
Overall 34.0 52.2 59.0 32.7 ($-1.3\%$) 53.3 ($+1.0\%$) 59.0 (+0.0%) 35.1 ($+1.1\%$) 53.5 ($+1.3\%$) 61.0 ($+2.0\%$)

### v2

@@ 4.1. Models
The selection of evaluated LLMs balances two practical factors: the ability to run an LLM locally and the availability of cloud endpoints with competitive token pricing. For the consumer-GPU category, we select GLM-4-9B-chat (glm2024chatglm) and Qwen2.5-Coder-7B-instruct (hui2024qwen2). Both are fully open-source, fit comfortably on a single 24 GB consumer GPU, and therefore incur no API costs. To represent low-cost hosted offerings, we add Qwen2.5-Plus (hui2024qwen2) and DeepSeek-V3 (liu2024deepseek). Qwen2.5-Plus is a closed-weight variant of the Qwen2.5 family (the provider does not disclose an exact parameter count) and is priced at \$0.11 per million input tokens and \$0.27 per million output tokens. DeepSeek-V3 is larger (671B parameters, 37B active during decoding) yet still affordable at \$0.27 per million input tokens and \$1.11 per million output tokens. All select LLMs are the standard chat versions rather than the more expensive thinking variants (such as DeepSeek-R1), ensuring a fair and comparable inference budget. This selection enables us to examine how input reduction behaves on both locally deployed LLMs and cost-efficient cloud LLMs.
@@ 4.2. Benchmark
A fair study of input reduction for automated program repair (APR) requires two features that existing datasets lack: (1) long failure-inducing test inputs and (2) low risk of training leakage. Widely used suites such as Defects4J (just2014defects4j), Human-EvalFix (muennighoffoctopack), and TutorCode (yang2024cref), Codeflaws (7965296), and the subset of LeetCode collected by Fan (fan2023automated) primarily contain only tiny tests, usually a few hundred characters, thus they do not reveal how well an APR pipeline copes when the failing input grows to tens of kilobytes. Meanwhile, most of those benchmarks were released years ago and are drawn from popular open-source projects that large language models have almost certainly seen, which can overstate repair accuracy. To our knowledge, no existing benchmark simultaneously offers long failure-inducing tests and low leakage risk; therefore, we need to build a new benchmark to fill this gap.
To meet the two requirements, the data source must publish every test file, including the largest hidden cases, and it must appear after the training cut-off dates of selected LLMs. LeetCode and Codeforces (including its variant Codeflaws) do not satisfy the first point because they do not share or share only small samples. AtCoder, by contrast, released the full test archives for every Beginner Contest (ABC) up to ABC 377, providing us with large inputs along with an official oracle. Therefore, we select all the satisfied tasks from ABC contests numbered 361 to 377, a span entirely after the knowledge cut-offs of the 4 LLMs we evaluate. From each contest, we retain tasks whose largest official test file is at least 4 KB and whose difficulty level is between C and F, ensuring that the tasks remain solvable for LLMs. To make input diversity explicit, we also categorize the input formats into 6 families and report coverage in Table 1. For cross-language validation, we additionally prepare LFTBench-Py, which mirrors the same 20 tasks and reuses the official tests; each task contributes one wrong-answer Python submission. For each task, we manually collect 10 C++ and 1 Python submissions that failed on a large test before July 1, 2025, resulting in 200 and 20 bugs, respectively. The median failing-input size is over 1 MB, and the largest single file exceeds 8 MB, which is large enough to trigger the “lost-in-the-middle” effect.
Table 1. Input-format categories with counts and tasks within LFTBench.
Category #Tasks Tasks
Single sequence 8 361C, 367D, 368C, 369D
370D, 373E, 377C, 377F
Multiple sequences 4 374E, 375C, 376C, 376E
Sequence with dependencies 4 364D, 366C, 372E, 371D
2-D matrix 1 363E
Graph 2 362D, 376D
String 1 365D
We also include a subset of OSS-Fuzz to cover repository-level defects. OSS-Fuzz continuously exercises project-specific fuzzers across many codebases and yields triaged, reproducible crash inputs with sanitizer stacks, which improves external validity and naturally exposes long and irregular inputs (googleAnnouncingOSSFuzzContinuous). Although some OSS-Fuzz repositories may overlap with LLM pretraining corpora, we hold the LLM and evaluation instances fixed across strategies, so relative differences between prompt types remain valid despite potential data leakage.
@@ 4.3. Metrics
A repair attempt succeeds when at least one candidate patch passes the full test suite (Eq. 3). For each bug $b$, we generate $n_{b}$ candidates once and observe $c_{b}$ passes. We report
$\text{pass@}k=\frac{1}{|\mathcal{B}|}\sum_{b\in\mathcal{B}}\left(1-\frac{\binom{n_{b}-c_{b}}{k_{b}}}{\binom{n_{b}}{k_{b}}}\right),\quad k_{b}=\min(k,n_{b}).$
This treats the $k$ samples as drawn without replacement and ignores ordering. We report pass@1, pass@5, and pass@10. These three cut-offs align with how developers inspect automated suggestions in practice. Kochhar et al. (kochhar2016practitioners) observe that most developers stop using a debugging tool if it does not help within the first five attempts, and Noller et al. (noller2022trust) find that few users review more than ten ranked patches. The same thresholds are widely adopted in recent APR work (huq2022review4repair; fu2022vulrepair; luo2025unlocking; yang2025morepair). Thus pass@1 reflects a one-shot setting, while pass@5 and pass@10 model realistic batch sizes that can be screened offline without taxing developer patience.
@@ 4.4. Hyperparameters
Table 2. Key hyper-parameters in ReduceFix.
@@
Reducer generation LLM backend Qwen2.5-Plus
@@
Wall-clock limit 60s per reduction
Repair inference # Samples per bug (pass@k) $k=\{1,5,10\}$
@@
To facilitate replication, the full artifact is published at
https://github.com/GLEAM-Lab/ReduceFix
@@ 5.1. Research Questions
  • RQ-1: How reliable are LLM-generated reducers at shrinking failure-inducing inputs? We assess the ReduceFix reduction phase through two metrics: the success rate and the compression ratio. Both metrics are evaluated against the unreduced original failure-inducing input, a ddmin-only approach, and a purely LLM-based input reduction approach.
  • RQ-2: Does supplying the reduced counterexample improve LLM-based repair? Using 4 LLMs, we test 3 prompting conditions: Baseline (no failure-inducing input), Origin Test (the full failure-inducing input), and Reduced Test (the reduced input produced by ReduceFix) on LFTBench, and report pass@k for $k\in\{1,5,10\}$. To assess cross-language portability, we also run ReduceFix on LFTBench-Py with Qwen2.5-Plus under the same settings.
@@
  • RQ-4: Can reduced-input prompting complement existing LLM-based APR pipelines? We integrate the ReduceFix into the ChatRepair (xia2024automated) and CREF (yang2024cref) without modifying its logic. Comparing the augmented version against the original one measures whether input reduction provides a drop-in boost for third-party APR systems.
  • RQ-5: How well does our approach perform on realistic repository-level repair scenarios? We evaluate ReduceFix end-to-end across OSS-Fuzz projects (googleAnnouncingOSSFuzzContinuous). For reduction, we compare ReduceFix against the ddmin-only and pure-LLM baselines. For repair, we compare pass@k across three prompt settings, the same as RQ-2, to assess whether reduced inputs improve repair accuracy.
@@ 5.2. RQ-1: Effectiveness of LLM-generated Reducer
[Experimental Design]: We assess the reduction performance on LFTBench, which contains 200 buggy C++ submissions drawn from 20 AtCoder Beginner Contest tasks. For each task, we prompt Qwen2.5-Plus once, using a one-shot example (ABC330D) and the problem statement of the task to produce reducer.py (see Section 3.2). The script then runs a ddmin loop on the full failure-inducing test input to find a smaller input that still causes the output of the buggy and reference programs to differ. To test whether ddmin is necessary, we add a baseline named pure-LLM, where the same LLM tries to generate a shorter test input directly. Furthermore, we introduce a baseline named ddmin-only, which directly applies the ddmin algorithm to the failure-inducing test input, using spaces and newlines as delimiters.
A reduction is counted as “Success” when it finishes on time without errors, and the failure is preserved, and returns a strictly smaller input.
Table 3. Reducer success and mean compression by difficulty. Compression is averaged over successful reductions.
Difficulty (n) Success Rate (%) Mean/Median Compression Rate (%)
ReduceFix Pure LLM ddmin-only ReduceFix Pure LLM ddmin-only
C (60) 100.0 63.3 30.0 84.5/100.0 100.0/100.0 100.0/100.0
D (80) 96.2 36.2 17.5 97.0/100.0 100.0/100.0 100.0/100.0
E&F (60) 88.3 21.7 65.0 83.0/99.9 99.1/99.2 84.2/99.9
Overall (200) 95.0 40.0 35.5 89.1/100.0 99.8/100.0 91.3/100.0
Table 4. Reducer success and mean compression by input format. Compression is averaged over successful reductions.
Input format (n) Success Rate (%) Mean/Median Compression Rate (%)
ReduceFix Pure LLM ddmin-only ReduceFix Pure LLM ddmin-only
Single sequence (80) 97.5 36.2 52.5 85.9/100.0 99.9/100.0 100.0/100.0
Multiple sequences
100.0 50.0 22.5 76.8/100.0 99.5/100.0 31.7/0.3
Sequence with
dependencies (40)
85.0 22.5 35.0 99.0/100.0 100.0/100.0 100.0/100.0
2-D matrix (10) 90.0 0.0 60.0 99.8/99.7 N/A 99.8/99.8
Graph (20) 95.0 65.0 0.0 100.0/100.0 100.0/100.0 N/A
String (10) 100.0 90.0 0.0 100.0/100.0 100.0/100.0 N/A
Overall (200) 95.0 40.0 35.5 89.1/100.0 99.8/100.0 91.3/100.0
@@
Method Pure LLM ReduceFix Ratio
@@
[Experimental Results]: Using the prompt template in Listing 1, Qwen2.5-Plus produced a syntactically valid reducer.py for all 200 buggy codes. Hence, all subsequent failures are due to the search phase rather than code-generation errors. Tables 3 and 4 show that the LLM-generated reducers succeed on 95.0% of the 200 buggy codes. All submissions of C-difficulty tasks are successfully reduced, while D-difficulty tasks and E & F-difficulty tasks reach 96.2% and 88.3% success, respectively. 10 of 200 forwarded the unreduced case to the repair phase, and in all 10 cases, the ddmin loop did not yield a shorter candidate before the time limit. The figure in README of our repository lists all 20 synthesized reducers, highlighting their structural decompositions and preserved semantic invariants.
Statistics in Tables 3 and 4 report the compression ratio, defined as Eq. 2, which means the percentage of bytes eliminated by the reducer. Under this definition, larger numbers indicate stronger compression: a value of 100% indicates that the input was reduced to almost nothing, while 0% means no reduction at all. The median removal rate remains near 100% across all difficulty levels, indicating that at least half of the test inputs that induce failures are nearly fully stripped yet still reproduce the bug. Averaged by difficulty, the reducer removes 83.0% of bytes on E & F-difficulty tasks, leaving 17.0% of the original data. Tasks in the C-difficulty and D-difficulty groups achieve even larger average removals of 84.5% and 97.0%, respectively.
Tables 3 and 4 show that directly conducting ddmin on the input achieves only a 35.5% success rate. Since input formats are often diverse, directly applying ddmin often results in a large number of invalid format attempts and thus exhibits a low reduction success rate. Tables 3 and 4 also show what happens when the LLM is leveraged to generate a shorter test input in a one-shot setting, without the pipeline of ReduceFix. This pure LLM baseline succeeds on only 40.0% of the bugs overall and 36.2% of the D-difficulty tasks, and by input type ReduceFix remains robust on graph and string cases where ddmin-only fails.
Table 5 compares the input/output token usage of two reduction strategies on LFTBench. The pure LLM approach, where the LLM tries to generate a shorter input in one pass, consumes about 5.7 million input tokens in total. The ReduceFix requires only 0.094 million input tokens for the same dataset. At current API prices, this comparison translates to \$0.632 versus \$0.017, 98% saving. Two factors drive this saving. First, the reducer script is generated on a per-task basis and then reused for every buggy submission of this task. Second, prompts in ReduceFix do not include the original failure-inducing input, which is often very large in LFTBench; they contain only the buggy code and the compact candidate input produced by ddmin. The pure LLM baseline must embed the entire long test case in the prompt, which greatly inflates the token cost.
[Failed Case Study]: For submission 62869553 of task ABC372E, the student program maintains, for every disjoint-set root, an array that stores the twenty largest vertex identifiers encountered so far. A Type 1 query merges two components by copying only the ten largest numbers from the other component, then sorts the twenty numbers in descending order. A Type 2 query asks for the $k$-th largest vertex inside the root that contains $v$. When a large component is merged into a smaller one, some very large identifiers disappear; later requests with $k\geq 9$ therefore return incorrect answers.
@@
We can resolve the issue by adding one simple guard. After the reducer renumbers the remaining vertices, it immediately reruns the interestingness test. If the failure no longer reproduces, the script rolls back to the original identifiers for that iteration instead of accepting the renumbered version. This small check, implemented in a few lines of Python, prevents the defect from being masked and allows the reduction process to continue correctly without any other modifications.
[RQ-1] [Findings]: (1) ReduceFix successfully reduced on 95% of the 200 bugs, significantly higher than Pure LLM and ddmin-only baselines. (2) Reductions delete 80%–97% of inputs on average, leaving a minimal yet still failing input. [Insights]: (1) These results validate the core design of ReduceFix: leveraging LLM to generate a reducer and then driving a systematic ddmin search is both essential and efficient. (2) The success of this “LLM + classic search” pattern suggests that similar hybrids could improve other code-analysis tasks.
@@ 5.3. RQ2: Effectiveness of ReduceFix
[Objective]: We determine whether steering the repair model with the counter-example reduced by ReduceFix improves the accuracy of generating a correct patch compared with (i) no failure-inducing input and (ii) the full failure-inducing input.
[Experimental Design]: We evaluate whether Reduced Test outperforms both Baseline and Origin Test across pass@k for multiple LLMs on LFTBench (C++), and whether this advantage holds on LFTBench-Py (Python) under the same settings. For every one of the 4 selected LLMs (Qwen2.5-Coder-7B-instruct, GLM-4-9B-chat, Qwen2.5-Plus, DeepSeek-V3), we test three prompting modes: Baseline (no test), Origin Test(full failure-inducing test), and Reduced Test. All other hyperparameters are fixed (temperature 0.8, $k\in\{1,5,10\}$, 10 candidate patches per bug), as shown in Table 2. Results are grouped by task difficulty. We also evaluate an end-to-end ddmin-only baseline that uses the reduced input when ddmin succeeds and otherwise falls back to the original failing test, with all other settings unchanged. In addition, for cross-language validation, we evaluate Qwen2.5-Plus on LFTBench-Py under the same protocol; the reducer succeeds on 18 of 20 Python cases (90%).
Table 6. Pass@K (%) across 4 LLMs.
Difficulty Baseline (No Test) Origin Test Reduced Test
@1 @5 @10 @1 @5 @10 @1 @5 @10
@@
C 5.5 17.1 23.3 3.7 13.2 20.0 5.5 16.7 25.0
D 6.4 17.4 25.0 6.1 17.4 23.8 9.9 26.0 36.2
E&F 1.3 5.9 10.0 1.8 7.1 11.7 2.3 8.5 11.7
Overall 4.6 13.9 20.0 4.1 13.1 19.0 6.3 17.9 25.5
GLM-4-9B-chat
C 2.8 5.8 8.3 1.3 3.8 5.0 2.2 4.1 5.0
D 3.5 10.6 13.8 2.0 7.3 11.2 5.8 14.5 20.0
E&F 0.7 1.6 1.7 0.5 1.5 1.7 1.2 1.7 1.7
Overall 2.5 6.5 8.5 1.3 4.5 6.5 3.3 7.5 10.0
Qwen2.5-Plus
C 37.3 60.3 68.3 31.8 53.5 63.3 33.7 53.5 65.0
D 40.1 55.6 60.0 39.9 58.7 61.3 43.6 58.7 62.5
E&F 22.5 39.7 48.3 24.0 45.7 51.7 25.0 46.5 55.0
Overall 34.0 52.2 59.0 32.7 53.3 59.0 35.1 53.5 61.0